# MDDM: A Molecular Dynamics Diffusion Model to Predict Particle Self-Assembly

### v2

@@ 5 Experiments
To evaluate our diffusion model, we pass a set of MD inputs into the model and compare the resulting structures using the radial distribution function (RDF). For the unconditional generation task, we see a close match between the generated outputs and the ground truth – this, and the low MSE value indicate that the denoising procedure can indeed perform conformation generation. The baseline, however, struggles to yield performance as strong as our model, likely due to it prioritizing recreation of global shape information rather than local structure.
@@
As for conditional generation, passing in potential as the input and using all 1000 MD results as training data (an additional 25 simulations for new parameters randomly sampled from the same range were used for testing), the results are promising, but leave room for improvement. Diffusion with Transolver still struggles for this task, but MDDM matches the first RDF peak well. Subsequent peaks, however, exhibit some mismatch, so there is still nuance to the conditional task that our model struggles to capture. To address this, starting with a smaller dataset, a larger model, or modified architecture (e.g. passing in the condition using cross-attention rather than concatenation), are a few strategies worth trying.
@@
Using a single NVIDIA RTX A6000 GPU with 128GB DDR4 RAM, performance metrics were evaluated for two models: Transolver and PBCConv. Transolver, with 128,867 parameters, achieved a training time of 24.03 seconds per 1,000 iterations and a sampling time of 2.53 seconds for 500 steps per data. In comparison, PBCConv, with 101,891 parameters, recorded a slightly faster training time of 23.15 seconds per 1,000 iterations but a longer sampling time of 4.17 seconds for 500 steps per data. The increased sampling time for our method is potentially attributed to the repeated graph generation, which requires calculating each particle’s relative distance to its neighbors at every step. This added computational complexity likely contributes to the observed difference in sampling efficiency.
@@ 6 Conclusion
We recontextualize diffusion models in a new domain: conformation generation for molecular dynamics material self-assembly. The proposed MDDM model uses a custom periodic graph neural network to iteratively denoise uniformly distributed particles into meaningful structures. For training data, we generated a dataset of 1000 MD results for different input potentials and temperatures. MDDM significantly outperforms point-cloud-based diffusion models for MD particle conformation generation. Unconditional generation has strong results, but there is still room for improvement in the conditional generation task.
For future work, we hope to improve conditional generation; some possible avenues toward achieving this may include increasing the capacity of the model or redesigning the model architecture. Exploration of MD systems with non-isotropic and/or many-body potential functions would be another interesting extension of this work.
@@ Acknowledgments
We thank the following individuals for their thoughtful discussion: Thomas O’Connor, Michael Bockstaller, Ayesha Abdullah, Andrew Gillman, Eric Harper, Jeffrey Ethier, Larry Drummy, Jacob Rast, and Matt Gormley.
Appendix Appendix A Figures
@@
Appendix Appendix B Shape generation task
@@
Appendix Appendix C Mathematical claims
Appendix C.1 Uniformity of a wrapped Gaussian
@@
Let $X$ be a random variable with a standard normal distribution. According to the Central Limit Theorem, this can be closely approximated by an Irwin-Hall distribution with $N=12$ (that is, the distribution describing the sum of 12 independent uniform random variables on $[0,1)$, subtracting 6 here to re-center at zero), which has mean 0 and standard deviation 1 when expressed using the PDF below, adapted from Marengo et al. [20].
@@
Suppose $Y=g_{L}(X)$ is a periodic boundary function of $X$, where $g_{L}(a)$ is defined such that $g_{L}(a)=a+k_{a}L$, for the unique integer $k_{a}$ such that $0\leq g_{L}(a)<L$. We aim to derive the PDF of $Y$. For a general non-monotonic transformation, the PDF of $Y$ is given by the change of variables:
@@
$f_{Y}(y)=\begin{cases}{\sum_{k=-6}^{5}\frac{1}{2(12-1)!}\sum_{r=0}^{12}(-1)^{r}\binom{12}{r}\text{sgn}(y+k+6-r)(y+k+6-r)^{12-1}}&\text{if }0\leq y<1,\\ 0&\text{otherwise}\end{cases}$
@@
$f_{Y}(y)=\begin{cases}1&\text{if }0\leq y<1,\\ 0&\text{otherwise}\end{cases}$
@@
Appendix C.2 Variational Inference for a Reinterpretation of Diffusion

### v3

@@ 3 Experiments and Results
Figure 5 shows a small set of conditional generation results on testing data. Here, the potential and temperature parameters were passed to the model as an additional set of node-wise input channels, now using all 1000 MD results as training data. Test data was generated by running an additional 25 simulations with new potential/temperature parameters uniformly randomly sampled from the training range. Diffusion with Transolver still struggles for this task, but MDDM matches the first RDF peak well. In an MD setting, the first peak is related to the coordination number of the RDF, which can correlate to the phase of the resulting material [26, 27]. Therefore, even this initial model may have utility, e.g. as a material phase classifier.
Subsequent peaks in the RDF, however, exhibit some mismatch, so there is still nuance to the conditional task that our model struggles to capture. To address this, updating the diffusion framework to condition results using a cross-attention or guidance-based method is worth investigating. For example, Classifier Guidance [28] and Classifier-Free Guidance [29] offer techniques for steering diffusion models toward high-quality conditioned results by “guiding” the denoising process. Methods like this are more sophisticated than the simple concatenation-based conditioning used in this work, and have greater potential to accurately model variations in potential and temperature.
@@ 4 Conclusion
We recontextualize diffusion models in a new domain: conformation generation for molecular dynamics material self-assembly. Through a redefinition of the forward noising process, our method enables potential-conditioned reverse diffusion of particles across periodic boundaries.
The proposed MDDM model uses a custom periodic graph neural network to iteratively denoise uniformly distributed particles into meaningful structures. For training data, we generated a dataset of 1000 MD self-assembly results for different input potentials and temperatures. The trained model takes potential/temperature information as a condition, and it outputs a corresponding thermodynamically valid particle conformation. MDDM significantly outperforms point-cloud-based diffusion models for MD particle conformation generation. Unconditional generation has strong results, but there is still room for improvement in the conditional generation task.
In future work, we hope to improve the results of conditional generation by leveraging alternative diffusion model conditioning methods. We also plan to investigate the effects of increasing the capacity of the model, as well as redesigning its architecture, or evaluating how the model can perform for MD materials systems with different underlying potentials. In particular, exploration of MD systems with non-isotropic and/or many-body potential functions would be another interesting extension of this work.
@@ Acknowledgments
This research was supported in part by the Center of Excellence established between the Air Force Research Laboratory and Carnegie Mellon University (AFOSR FA8650-19-5209). We thank the following individuals for their thoughtful discussion: Thomas O’Connor, Michael Bockstaller, Ayesha Abdullah, Andrew Gillman, Eric Harper, Jeffrey Ethier, Larry Drummy, Jacob Rast, and Matt Gormley.
Appendix Appendix A Dataset visualization
@@
Appendix Appendix B Shape generation demonstration
@@
Appendix Appendix C Code listings
This section contains Python code for simplified versions of the modules used for the denoiser in MDDM. PeriodicBox() is a utility for handling periodic boundary conditions, while PBCNet() is a module that contains several PBCConv layers, as defined in Fig. 3 and Eqn. 1. Each of these are showcased below.
Appendix C.1 PeriodicBox()
This module provides several utilities for applying periodic boundaries, which is necessary for dealing with systems of particles. It contains the method wrap_nearest() for computing the shortest vector from one particle to another, wrapping across the periodic boundaries as necessary. wrap_within() wraps a particle’s position back within the periodic box, for example if noising had removed it. Methods knn_tensor() and knn_edges() enable computation of a k-nearest neighbor graph under the presence of PBCs. The PeriodicBox() module is used extensively throughout the diffusion model, to wrap across periodic boundaries and to determine nearest neighbors in the periodic domain.
Listing 1: PeriodicBox module, for handling wrapping around periodic boundaries, including creating of k-nearest neighbors graph
class PeriodicBox($n$.Module):
def __init__(self, Ls = torch.tensor([1., 1., 1.])):
super().__init__()
if type(Ls) != torch.tensor:
Ls = torch.tensor(Ls, dtype=torch.float)
self.L = Ls
def wrap_nearest(self, A, B): # Get vector from A to B, wrapping as necessary
dr = B - A
return dr + self.L*(dr < -self.L/2) - self.L*(dr > self.L/2)
def wrap_within(self, X): # Wrap points X to be within the periodic box
return X - self.L * (X > self.L) + self.L * (X < 0)
def wrap_dist(self, A, B): # Get distance from A to B, wrapping as necessary
return torch.norm(self.wrap_nearest(A, B), dim = -1)
def knn_tensor(self, pts, k):
D = self.wrap_dist(pts[:, None, :], pts[None, :, :])
nearest = torch.topk(D, k+1, dim=1, largest=False)[1][:,1:]
return nearest
def knn_edges(self, pts, k):
nearest = self.knn_tensor(pts, k)
src = torch.arange(nearest.size(0)).repeat_interleave(nearest.size(1))
dest = nearest.flatten()
edges = torch.stack((src, dest), dim=1).long()
return edges
Appendix C.2 PBCNet()
PBCNet() is a PyTorch module that serves as the denoising model in MDDM. The forward pass takes particle coordinates as input, computes the k-nearest neighbor graph, and then performs several periodic graph convolutions. Finally, each particle’s convolved features are passed through an MLP to compute an estimate of what noise was added to the data to arrive at the input. See Fig. 3 for the network structure and Eqn. 1 for the graph convolution formula.
Listing 2: PBCNet module, a graph neural network architecture in which nodes exchange absolute-position-invariant information with nearest neighbors, including across periodic boundaries
class PBCNet($n$.Module):
def __init__(self, n_dims, n_features, n_conv_features, n_out,
L=10, k=12, n_conv_layers=2,
conv_mlp_size=[64,64], out_mlp_size=[128,128],
periodic_box=None, residual=True, concat_features=True):
super().__init__()
self.n_dims, self.n_features = n_dims, n_features
self.n_conv_features, self.n_out = n_conv_features, n_out
self.k, self.n_conv_layers = k, n_conv_layers
self.residual, self.concat_features = residual, concat_features
self.pbc = periodic_box or PeriodicBox([L]*n_dims)
self.in_conv = MLP([n_dims+n_features, *conv_mlp_size, n_conv_features])
self.convs = $n$.ModuleList([
MLP([n_dims + 2*n_conv_features + (n_features if concat_features else 0),
*conv_mlp_size, n_conv_features])
for _ in range(n_conv_layers)
self.out_mlp = MLP([n_conv_features + (n_features if concat_features else 0),
*out_mlp_size, n_out])
def get_edges(self, pts):
"""Return neighbor edges and wrapped displacement vectors."""
edges = self.pbc.knn_edges(pts, self.k)
vecs = self.pbc.wrap_nearest(pts[edges[:,0]], pts[edges[:,1]])
return edges, vecs
def conv(self, mlp, vecs, edges, node_f=None, shared_f=None):
"""Apply one convolution layer and pool edge features."""
feature_vecs = [vecs]
if node_f is not None:
src, dst = node_f[edges[:,0]], node_f[edges[:,1]]
feature_vecs += [src, src - dst]
if shared_f is not None:
feature_vecs.append(shared_f[edges[:,0]])
A = torch.cat(feats, -1).float()
B = mlp(A).view(node_f.shape[0], self.k, self.n_conv_features)
return torch.max(B, 1)[0]
def forward(self, X):
# This implementation does not allow a batch dimension
coords, shared = X[:,:self.n_dims], X[:,self.n_dims:]
edges, vecs = self.get_edges(coords)
# Initial PBC-Conv layer
f = self.conv(self.in_conv, vecs, edges, shared_f=shared)
shared_f = shared if self.concat_features else None
# Subsequent PBC-Conv layers, optionally residual
for mlp in self.convs:
f_new = self.conv(mlp, vecs, edges, node_f=f, shared_f=shared_f)
f = f + f_new if self.residual else f_new
# Output particle-wise MLP
if self.concat_features: f = torch.cat([shared_f, f], -1)
return self.out_mlp(f)
Appendix C.3 MD Input Script
To generate a large dataset of MD results, we define the following string in a Python script. This string’s bracketed values are replaced using .format() and the result is written to a LAMMPS [18] input script file. 1000 such scripts are generated and simulated in LAMMPS to generate the dataset of bulk conformations. Each simulation anneals to the target temperature from 10 times the (absolute) target temperature over the course of 100,000 time steps in NVT, and then allows further equilibration for an additional 100,000 time steps. The dimensionless time step is set as 0.005. Particles are initialized in a 10x10x10 simple cubic lattice, perturbed slightly with uniform noise, and assigned random velocity vectors sampled from a standard normal distribution.
Listing 3: LAMMPS input script for running a 1000-particle MD self-assembly with a custom tabulated potential
IN_CUSTOM = """
units lj
atom_style atomic
dimension 3
boundary p p p
region box block 0 10 0 10 0 10
create_box 1 box
lattice sc 1.0
create_atoms 1 box
pair_style table linear 1000
pair_coeff 1 1 custom.table CUSTOM
mass 1 1.0
velocity all create 1.0 12345 dist gaussian
displace_atoms all random 0.1 0.1 0.1 12345
neighbor 0.3 bin
neigh_modify delay 0 every 1 check yes
# Nose-Hoover Thermostat for NVT
fix 1 all nvt temp {Ti} {Tf} \$(100.0*dt)
timestep 0.005
thermo 100
thermo_style custom step temp pe ke etotal
run {num_steps}
dump 1 all atom {dump_every} {dump_file}
run {num_steps}
Appendix Appendix D Mathematical claims
Appendix D.1 Uniformity of a wrapped Gaussian
@@
Let $X$ be a random variable with an approximately standard normal distribution. According to the Central Limit Theorem, a normal distribution can be closely approximated by an Irwin-Hall distribution with $N=12$ (that is, the distribution describing the sum of 12 independent uniform random variables on $[0,1)$, subtracting 6 in this case to re-center at zero), which has mean 0 and standard deviation 1 when expressed using the PDF below, adapted from Marengo et al. [30]:
@@
Suppose $Y=g_{L}(X)$ is a periodic boundary function of $X$, where $g_{L}(a)$ is defined as $g_{L}(a)=a+k_{a}L$, for the unique integer $k_{a}$ such that $0\leq g_{L}(a)<L$. We aim to derive the PDF of $Y$. For a general non-monotonic transformation, the PDF of $Y$ is given by the change of variables:
@@
$f_{Y}(y)=\begin{cases}{\sum_{k=-6}^{5}\frac{1}{2(12-1)!}\sum_{r=0}^{12}(-1)^{r}\binom{12}{r}\text{sgn}(y+k+6-r)(y+k+6-r)^{12-1}}&\text{if }0\leq y<1,\\ 0&\text{otherwise.}\end{cases}$
@@
$f_{Y}(y)=\begin{cases}1&\text{if }0\leq y<1,\\ 0&\text{otherwise.}\end{cases}$
@@
Appendix D.2 Variational Inference for a reinterpretation of forward diffusion